# latexText page

### v1

@@ Acknowledgements
Bottom-up GARCH (GARCH.BU): In this fully parametric model we assume a copula-GARCH model with skew-$t$ residuals on $\{\bm{X}_{t}\}_{t\in\mathbb{T}}$. For $j=1,\dots,d$, we assume an AR-GARCH(1,1) model $X_{j,t}=\mu_{j}+\sigma_{j,t}Z_{j,t}$ on the marginal time series $\{X_{j,t}\}_{t\in\mathbb{T}}$, where $\mu_{j}\in\mathbb{R}$ is a location parameter, and $\sigma_{j,t}>0$ is a conditional standard deviation following the GARCH(1,1) dynamics. The residual series $\{Z_{j,t}\}_{t\in\mathbb{T}}$ is assumed to be a strict white noise of a skew-$t$ distribution $\operatorname{Skt}(\nu_{j},\gamma_{j})$ (Fernández and Steel, 1998) for degrees of freedom $\nu_{j}>0$ and skewness parameter $\gamma_{j}>0$. For the dependence structure, we assume that $(Z_{1,t},\dots,Z_{d,t})$, $t\in\mathbb{T}$, has a $t$-copula with degrees of freedom $\nu>0$ and correlation matrix $P$. Under this model assumption, the joint distribution of $\bm{X}_{n+1}|{\cal G}_{n}$ is specified by the skew-$t$ margins and the $t$-copula. We first estimate parameters of the marginal GARCH models by the maximum likelihood estimation using the package fGarch (Wuertz et al., 2022) implemented in R (R Core Team, 2023). We then estimate parameters of the $t$-copula from $(Z_{1,t},\dots,Z_{d,t})$, $t\in\mathbb{T}_{\text{in}}$, by the maximum pseudo-likelihood estimation with inversion of Kendall’s tau (Demarta and McNeil, 2005), which is implemented in the pacakge copula (Hofert et al., 2023). Based on parameters’ estimates we forecast total VaR, total ES, and ES contributions of $\bm{X}_{n+1}|{\cal G}_{n}$ by Monte Carlo simulation with sample size $10^{5}$.
Top-down GARCH (GARCH.TD): In this model we assume a bivariate copula-GARCH model with skew-$t$ residuals on $\{(X_{j,t},S_{t})\}_{t\in\mathbb{T}}$ for each $j\in\{1,\dots,d\}$, where the marginal AR-GARCH model on $\{S_{t}\}_{t\in\mathbb{T}}$ is shared in common. As in the GARCH.BU model, these assumptions enable the prediction of total VaR, total ES, and ES contributions of $\bm{X}_{n+1}|{\cal G}_{n}$ through Monte Carlo simulation. Since we fit a common AR-GARCH(1,1) model $S_{t}=\mu+\sigma_{t}Z_{t}$ on the series $\{S_{t}\}_{t\in\mathbb{T}}$ and directly model the copula of $(Z_{j,t},Z_{t})$, $t\in\mathbb{T}$, for $j=1,\dots,d$, the model of the vector of residuals $(Z_{1,t},\dots,Z_{d,t})$, $t\in\mathbb{T}$, is left unspecified.
Hysteretic autoregressive model with GARCH error and dynamic conditional correlations (HAR.GARCH): In this model we consider the bivariate hysteretic autoregressive (HAR) model with GARCH error and dynamic conditional correlations (Tse and Tsui, 2002; Chen et al., 2019) on the time series $\{(X_{j,t},S_{t})\}_{t\in\mathbb{T}}$ for $j=1,\dots,d$. We say that a pair of asset return series $\bm{y}_{t}=(y_{1,t},y_{2,t})$ follows an HAR.GARCH model if:
@@
Parameters of the HAR.GARCH model include $\bm{\Phi}_{0}^{(k)}\in\mathbb{R}^{2}$, $\bm{\Phi}_{1}^{(k)}\in\mathbb{R}^{2\times 2}$, $r_{L},r_{U}\in\mathbb{R}$, and a positive-definite matrix $\bm{\Gamma}^{(k)}\in\mathbb{R}^{2\times 2}$ with unit diagonal elements and non-negative real numbers $\alpha_{i,0}^{(k)}$ ,$\alpha_{i,1}^{(k)}$, $\beta_{i,1}^{(k)}$, $\theta_{1}^{(k)}$, and $\theta_{2}^{(k)}$ such that $10<\alpha_{i,1}^{(k)}+\beta_{i,1}^{(k)}<$ 1 and $10<\theta_{1}^{(k)}+\theta_{2}^{(k)}<$ 1 for $i,k\in\{1,2\}$. For a hysteresis variable, we utilize an endogenous variable $z_{t}=y_{1,t}$ to exhibit hysteresis effects. Following Chen et al. (2019), we choose an adapted multivariate Student t distribution, based on the scale mixtures of the normal representation, as the distribution $D(0,\bm{\Gamma}_{t})$ for the bivariate vector of residuals $\bm{\epsilon}_{t}$. We utilize Bayesian methods to estimate parameters of the above model and the risk quantities of interest; see Sections 3 and 4 of Chen et al. (2019), respectively. In our setting, $y_{1,t}$ stands for $X_{j,t}$, $j\in\{1,\dots,d\}$, and $y_{2,t}$ represents $S_{t}$. As the total VaR is predicted $d$-times, we simply use the average of them. Moreover, total ES is predicted as the sum of ES contributions.
Compositional regression model with least square estimation (CR.LSE): Following Boonen et al. (2019), we first obtain a series of ES contributions by the elliptical formula (Corollary 8.43 of McNeil et al., 2015):
$\widehat{ESC}_{t}^{\text{EL}}$ $=\left(\widehat{ESC}_{1,t}^{\text{EL}},\dots,\widehat{ESC}_{d,t}^{\text{EL}}\right)$
@@
where $\hat{\bm{\mu}}=(\hat{\mu_{1}},\dots,\hat{\mu_{d}})$ is the vector of location parameters, and $\hat{\Sigma}_{t}$ is the conditional covariance matrix estimated in GARCH.BU. We also estimate the dynamics of total VaR and total ES parametrically based on the AR-GARCH(1,1) assumption on $\{S_{t}\}_{t\in\mathbb{T}}$ with skew-$t$ residuals. We then obtain the allocation weights as $\hat{\bm{w}}_{t}=\operatorname{C}\left(\widehat{ESC}_{t}^{\text{EL}}\right)$, $t\in\mathbb{T}$. Regarding this set of allocation weights as compositional data, we estimate the parameter $\bm{\theta}$ of $\upsilon_{\bm{\theta}}$ in  (9) as follows:
$\hat{\bm{\theta}}=\operatorname{argmin}_{\bm{\theta}\in\Theta}\sum_{t=1}^{n-1}\sum_{j=1}^{d}\left\{\hat{w}_{j,t+1}-\upsilon_{\bm{\theta}}(\hat{\bm{w}}_{t},\bm{X}_{t})_{j}\right\}^{2}.$
Finally, we use this $\hat{\bm{\theta}}$ to predict ES contributions of $\bm{X}_{n+1}|\mathcal{G}_{n}$ following Section 4.
Compositional regression model based on score minimization (CR.OPT): This is the proposed model described in Section 4. For the dynamics of total VaR and total ES, we use the same estimates used in CR.LSE. We choose $\bm{w}_{1}=\operatorname{C}\left(\widehat{ESC}_{1}^{\text{EL}}\right)$ for the initial allocation weight.
S2.3 Time series plots
Figures S2 and S3 display the estimated dynamics of total VaR, total ES, and ES contributions on the out-of-sample period $\mathbb{T}_{\text{out}}$ together with the realized losses. In Figure S2 we do not observe a major difference among the estimates of total VaR and total ES except HS, which only captures the average trends. This feature of HS is also observable for ES contributions in Figure S3. For the estimates of ES contributions, other distinguishable models are HAR.GARCH and CR.OPT, where the former tends to estimate ES contributions lower than others, and the latter typically leads to more fluctuated estimates. It may be worth noting that the average score (10) minimized in CR.OPT does not take non-extreme losses into account, and thus the estimates of CR.OPT can deviate from those of other models in normal times. Compared with HS, HAR.GARCH, and CR.OPT, the other three models seem to produce estimates relatively similar with each other.
Figure S2: Time series plots of total VaR and total ES with confidence level $0.975$ estimated by the $6$ models. The black lines represent the time series of the total loss of the portfolio.
Figure S3: Time series plots of ES contributions with confidence level $0.975$ estimated by the $6$ models. The black lines represent the time series of the corresponding marginal losses.
S2.4 Backtests in the one-step approach
Due to multi-objective elicitability of ES contributions, we also conduct two-sided and “one and a half-sided” Wald-tests for tuple of ES contributions and each of them in combination with total VaR; see Section 5 of Fissler and Hoga (2023) for details of these tests. In this one-step approach the major difference from the two-step approach is that the null hypothesis includes equality in the accuracy of total VaRs for two competing forecasts . We report p-values of these tests in Table S5.
Table S5: Results of the Wald-tests to compare the forecast accuracy of ES contributions with CR.OPT as the benchmark model.
[…]
@@
(1) Tuple of ES contributions (2) ES contribution (AMZN)
HS 0.007 0.012 0.004 HS 0.010 0.012 0.006
GARCH.BU 0.126 0.368 0.084 GARCH.BU 0.466 0.368 0.342
GARCH.TD 0.315 0.320 0.222 GARCH.TD 0.288 0.320 0.201
HAR.GARCH 0.316 0.929 0.223 HAR.GARCH 0.830 0.685 0.929
CR.LSE 0.735 1.000 0.584 CR.LSE 0.478 1.000 0.351
@@
(3) ES contribution (GOOGL) (4) ES contribution (TSLA)
HS 0.019 0.012 0.012 HS 0.011 0.012 0.007
GARCH.BU 0.453 0.368 0.330 GARCH.BU 0.098 0.368 0.065
GARCH.TD 0.406 0.292 0.320 GARCH.TD 0.359 0.320 0.255
HAR.GARCH 0.656 0.929 0.507 HAR.GARCH 0.328 0.929 0.231
CR.LSE 0.742 0.591 1.000 CR.LSE 0.788 1.000 0.639
@@
  • ${}^{a}$he p-values are calculated based on the three different null hypotheses with all of them including the equality of the accuracy of forecasted total VaRs, where “$=$ CR.OPT” means that the model is equally accurate as CR.OPT, and $\leq$ ($\geq$) CR.OPT” represents the hypothesis that the model is less (more) accurate than CR.OPT.
Due to joint identifiability of total ES (see Section 2.1 of  Nolde and Ziegel, 2017) and $j$th ES contribution for $j=1,\dots,d$ (see Proposition 2) in combination with total VaR, we also conduct Wald-tests for these risk quantities in the one-step approach. Namely, for total ES, we consider the two-sided Wald-test for the null hypothesis:
@@
based on Section 5 of Fissler and Hoga (2023) as considered in the above comparative backtests. For the $j$th ES contribution for $j\in\{1,\dots,d\}$, we replace $\bar{\operatorname{V}}_{T}^{\text{ES}}=(1/T)\sum_{t=1}^{T}\operatorname{V}_{t}^{\text{ES}}$ above with $\bar{\operatorname{V}}_{j,T}^{\text{ESC}}=(1/T)\sum_{t=1}^{T}\operatorname{V}_{j,t}^{\text{ESC}}$. We consider these null hypotheses since over- and under-estimations of ES contributions can be judged when the corresponding total VaR is calibrated; see Section 2.3. We summarize p-values of these tests in Table S6.
Table S6: Results of the Wald-tests to verify forecast accuracy of total ES and ES contributions.
[…]
@@
GARCH.BU 0.014 0.009 0.034 GARCH.BU 0.052 0.034 0.034
GARCH.TD 0.061 0.040 0.204 GARCH.TD 0.266 0.185 0.204
@@
CR.LSE 0.174 0.118 0.204 CR.LSE 0.263 0.183 0.204
CR.OPT 0.174 0.118 0.204 CR.OPT 0.232 0.159 0.204
(3) ES contribution (GOOGL) (4) ES contribution (TSLA)
@@
GARCH.BU 0.052 0.034 0.034 GARCH.BU 0.018 0.011 0.034
GARCH.TD 0.245 0.169 0.204 GARCH.TD 0.232 0.160 0.204
@@
CR.LSE 0.253 0.175 0.204 CR.LSE 0.274 0.191 0.204
CR.OPT 0.291 0.204 0.203 CR.OPT 0.251 0.173 0.204
  • ${}^{a}$he p-values are calculated based on the three different null hypotheses with all of them including the preciseness of forecasted total VaRs, where “$=$ True” means that the forecast is precise, $\leq$ True” represents the hypothesis that the forecast is under-estimated, and $\geq$ True” stands for the case when the forecast is over-estimated.

### v2

@@ Acknowledgements
Bottom-up GARCH (GARCH.BU): In this fully parametric model we assume a copula-GARCH model with skew-$t$ residuals on $\{\bm{X}_{t}\}_{t\in\mathbb{T}}$. For $j=1,\dots,d$, we assume an AR-GARCH(1,1) model $X_{j,t}=\mu_{j}+\sigma_{j,t}Z_{j,t}$ on the marginal time series $\{X_{j,t}\}_{t\in\mathbb{T}}$, where $\mu_{j}\in\mathbb{R}$ is a location parameter, and $\sigma_{j,t}>0$ is a conditional standard deviation following the GARCH(1,1) dynamics. The residual series $\{Z_{j,t}\}_{t\in\mathbb{T}}$ is assumed to be a strict white noise of a skew-$t$ distribution $\operatorname{Skt}(\nu_{j},\gamma_{j})$ (Fernández and Steel 1998) for degrees of freedom $\nu_{j}>0$ and skewness parameter $\gamma_{j}>0$. For the dependence structure, we assume that $(Z_{1,t},\dots,Z_{d,t})$, $t\in\mathbb{T}$, has a $t$-copula with degrees of freedom $\nu>0$ and correlation matrix $P$. Under this model assumption, the joint distribution of $\bm{X}_{n+1}|{\cal G}_{n}$ is specified by the skew-$t$ margins and the $t$-copula. We first estimate parameters of the marginal GARCH models by the maximum likelihood estimation using the package fGarch (Wuertz et al. 2022) implemented in R (R Core Team 2023). We then estimate parameters of the $t$-copula from $(Z_{1,t},\dots,Z_{d,t})$, $t\in\mathbb{T}_{\text{in}}$, by the maximum pseudo-likelihood estimation with inversion of Kendall’s tau (Demarta and McNeil 2005), which is implemented in the pacakge copula (Hofert et al. 2023). Based on parameters’ estimates we forecast total VaR, total ES, and ESCs of $\bm{X}_{n+1}|{\cal G}_{n}$ by Monte Carlo simulation with sample size $10^{5}$.
Top-down GARCH (GARCH.TD): In this model we assume a bivariate copula-GARCH model with skew-$t$ residuals on $\{(X_{j,t},S_{t})\}_{t\in\mathbb{T}}$ for each $j\in\{1,\dots,d\}$, where the marginal AR-GARCH model on $\{S_{t}\}_{t\in\mathbb{T}}$ is shared in common. Based on the AR-GARCH(1,1) assumption on $\{S_{t}\}_{t\in\mathbb{T}}$ with skew-$t$ residuals, we estimate the dynamics of total VaR and total ES parametrically. We then predict ESCs of $\bm{X}_{n+1}|{\cal G}_{n}$ through Monte Carlo simulation. Note that the model of the vector of residuals $(Z_{1,t},\dots,Z_{d,t})$, $t\in\mathbb{T}$, is left unspecified since we fit a common AR-GARCH(1,1) model $S_{t}=\mu+\sigma_{t}Z_{t}$ on the series $\{S_{t}\}_{t\in\mathbb{T}}$ and directly model the copula of $(Z_{j,t},Z_{t})$, $t\in\mathbb{T}$, for $j=1,\dots,d$.
Hysteretic autoregressive model with GARCH error and dynamic conditional correlations (HAR.GARCH): In this model we consider the bivariate hysteretic autoregressive (HAR) model with GARCH error and dynamic conditional correlations (Tse and Tsui 2002; Chen et al. 2019) on the time series $\{(X_{j,t},S_{t})\}_{t\in\mathbb{T}}$ for $j=1,\dots,d$. We say that a pair of asset return series $\bm{y}_{t}=(y_{1,t},y_{2,t})$ follows an HAR.GARCH model if:
@@
Parameters of the HAR.GARCH model include $\bm{\Phi}_{0}^{(k)}\in\mathbb{R}^{2}$, $\bm{\Phi}_{1}^{(k)}\in\mathbb{R}^{2\times 2}$, $r_{L},r_{U}\in\mathbb{R}$, and a positive-definite matrix $\bm{\Gamma}^{(k)}\in\mathbb{R}^{2\times 2}$ with unit diagonal elements and non-negative real numbers $\alpha_{i,0}^{(k)}$, $\alpha_{i,1}^{(k)}$, $\beta_{i,1}^{(k)}$, $\theta_{1}^{(k)}$, and $\theta_{2}^{(k)}$ such that $10<\alpha_{i,1}^{(k)}+\beta_{i,1}^{(k)}<$ 1 and $10<\theta_{1}^{(k)}+\theta_{2}^{(k)}<$ 1 for $i,k\in\{1,2\}$. For a hysteresis variable, we utilize an endogenous variable $z_{t}=y_{1,t}$ to exhibit hysteresis effects. Following Chen et al. (2019), we choose an adapted multivariate Student t distribution, based on the scale mixtures of the normal representation, as the distribution $D(0,\bm{\Gamma}_{t})$ for the bivariate vector of residuals $\bm{\epsilon}_{t}$. We utilize Bayesian methods to estimate parameters of the above model and the risk quantities of interest; see Sections 3 and 4 of Chen et al. (2019), respectively. In our setting, $y_{1,t}$ stands for $X_{j,t}$, $j\in\{1,\dots,d\}$, and $y_{2,t}$ represents $S_{t}$. As the total VaR is predicted $d$-times, we simply use the average of them. Moreover, total ES is predicted as the sum of ESCs.
Compositional regression model with least square estimation (CR.LSE): Following Boonen et al. (2019), we first obtain a series of ESCs by the elliptical formula (Corollary 8.43 of McNeil et al. 2015):
$\widehat{ESC}_{t}^{\text{EL}}$ $=\left(\widehat{ESC}_{1,t}^{\text{EL}},\dots,\widehat{ESC}_{d,t}^{\text{EL}}\right)^{\top}$
@@
where $\hat{\bm{\mu}}=(\hat{\mu_{1}},\dots,\hat{\mu_{d}})$ is the vector of location parameters, and $\hat{\Sigma}_{t}$ is the conditional covariance matrix estimated in GARCH.BU. For the dynamics of total VaR and total ES, we use the same estimates used in GARCH.TD. We then obtain the allocation weights as $\hat{\bm{w}}_{t}=\operatorname{C}\left(\widehat{ESC}_{t}^{\text{EL}}\right)$, $t\in\mathbb{T}$. Regarding this set of allocation weights as compositional data, we estimate the parameter $\bm{\theta}$ of $\upsilon_{\bm{\theta}}$ in  (9) as follows:
$\hat{\bm{\theta}}=\operatorname{argmin}_{\bm{\theta}\in\Theta}\sum_{t=1}^{n-1}\sum_{j=1}^{d}\left\{\hat{w}_{j,t+1}-\upsilon_{\bm{\theta}}(\hat{\bm{w}}_{s},\bm{X}_{s},s\leq t)_{j}\right\}^{2}.$
Finally, we use this $\hat{\bm{\theta}}$ to predict ESCs of $\bm{X}_{n+1}|\mathcal{G}_{n}$ following Section 4.
Compositional regression model based on score optimization (CR.OPT): This is the proposed model described in Section 4. For the dynamics of total VaR and total ES, we use the same estimates used in CR.LSE. We choose $\bm{w}_{1}=\operatorname{C}\left(\widehat{ESC}_{1}^{\text{EL}}\right)$ for the initial allocation weight. When optimizing (10), we use the vector of parameters estimated in CR.LSE as the initial value.
S2.3 Backtests in the one-step approach
Due to multi-objective elicitability of ESCs, we also conduct two-sided and “one and a half-sided” Wald-tests for tuple of ESCs and each of them in combination with total VaR; see Section 5 of Fissler and Hoga (2024) for details of these tests. In this one-step approach the major difference from the two-step approach is that the null hypothesis includes equality in the accuracy of total VaRs for two competing forecasts . We report p-values of these tests in Table S6.
Table S6: Results of the Wald-tests to compare the forecast accuracy of ESCs with CR.OPT as the benchmark model.
p-valuea p-valuea
@@
(1) Tuple of ESCs (2) ESC (AMZN)
HS 0.006 0.012 0.004 HS 0.010 0.012 0.006
GARCH.BU 0.054 0.404 0.035 GARCH.BU 0.533 0.398 0.404
GARCH.TD 0.524 1.000 0.390 GARCH.TD 0.963 1.000 0.874
HAR.GARCH 0.418 0.929 0.302 HAR.GARCH 0.739 0.588 0.929
CR.LSE 0.228 1.000 0.157 CR.LSE 0.358 1.000 0.255
@@
(3) ESC (GOOGL) (4) ESC (TSLA)
HS 0.019 0.012 0.012 HS 0.009 0.012 0.006
GARCH.BU 0.434 0.404 0.315 GARCH.BU 0.034 0.404 0.021
GARCH.TD 0.999 1.000 0.981 GARCH.TD 0.526 1.000 0.392
HAR.GARCH 0.559 0.929 0.420 HAR.GARCH 0.303 0.929 0.213
CR.LSE 0.907 0.782 1.000 CR.LSE 0.211 1.000 0.144
@@
  • aThe p-values are calculated based on the three different null hypotheses with all of them including the equality of the accuracy of forecasted total VaRs, where “$=$ CR.OPT” means that the model is equally accurate as CR.OPT, and $\leq$ ($\geq$) CR.OPT” represents the hypothesis that the model is less (more) accurate than CR.OPT.
Due to joint identifiability of total ES (see Section 2.1 of  Nolde and Ziegel 2017) and $j$th ESC for $j=1,\dots,d$ (see Proposition 2) in combination with total VaR, we also conduct Wald-tests for these risk quantities in the one-step approach. Namely, for total ES, we consider the two-sided Wald-test for the null hypothesis:
@@
based on Section 5 of Fissler and Hoga (2024) as considered in the above comparative backtests. For the $j$th ESC for $j\in\{1,\dots,d\}$, we replace $\bar{\operatorname{V}}_{T}^{\text{ES}}=(1/T)\sum_{t=1}^{T}\operatorname{V}_{t}^{\text{ES}}$ above with $\bar{\operatorname{V}}_{j,T}^{\text{ESC}}=(1/T)\sum_{t=1}^{T}\operatorname{V}_{j,t}^{\text{ESC}}$. We consider these null hypotheses since over- and under-estimations of ESCs can be judged when the corresponding total VaR is calibrated; see Section S1.1. We summarize p-values of these tests in Table S7.
Table S7: Results of the Wald-tests to verify forecast accuracy of total ES and ESCs.
p-valuea p-valuea
@@
GARCH.BU 0.008 0.005 0.025 GARCH.BU 0.038 0.024 0.025
GARCH.TD 0.174 0.118 0.204 GARCH.TD 0.270 0.188 0.204
@@
CR.LSE 0.174 0.118 0.204 CR.LSE 0.276 0.192 0.204
CR.OPT 0.174 0.118 0.204 CR.OPT 0.215 0.147 0.204
(3) ESC (GOOGL) (4) ESC (TSLA)
@@
GARCH.BU 0.037 0.024 0.025 GARCH.BU 0.011 0.007 0.025
GARCH.TD 0.238 0.164 0.204 GARCH.TD 0.230 0.158 0.204
@@
CR.LSE 0.264 0.184 0.204 CR.LSE 0.262 0.182 0.204
CR.OPT 0.275 0.191 0.204 CR.OPT 0.280 0.195 0.204
  • aThe p-values are calculated based on the three different null hypotheses with all of them including the preciseness of forecasted total VaRs, where “$=$ True” means that the forecast is precise, $\leq$ True” represents the hypothesis that the forecast is under-estimated, and $\geq$ True” stands for the case when the forecast is over-estimated.